OpenStudy (anonymous):

How can I look at an integral and know whether to use FTC or U-substitution?

zepdrix (zepdrix):

Do you mean FTC, Part 1?\[\Large\rm \frac{d}{dx}\int\limits_0^x f(t)dt=f(x)\]This one?

OpenStudy (anonymous):

Part one and two.

zepdrix (zepdrix):

You're always applying part two when you have `bounds on your integral`. You're apply the FTC, Part 1, when you're taking a derivative of an integral.

zepdrix (zepdrix):

I'm nooooooot exactly sure how that relates to U-substitution though hehe

OpenStudy (anonymous):

The thing is, I'm getting questions that ask me to evaluate the integral, so I'm not always sure which method to use because I think it includes u-substitution, right?

OpenStudy (jhannybean):

You use a u-substitution to make your integral easier to solve, therefore you can use Part 2 of FTC.

OpenStudy (jhannybean):

One is a theorem of integration - FTC , and one is a method of integration - u-substitution.

OpenStudy (anonymous):

oh okay!

OpenStudy (jhannybean):

When you're finding the integral of some function, you're basically finding the area the solid, or even function makes within a certain limit, a certain space. The Fundamental Theorem of Calculus tells you that from a starting point A to an ending point B, you will be integrating so and so function to find area, volume, etc.

OpenStudy (anonymous):

thank you both so much!

OpenStudy (jhannybean):

The FTC holds for all integrals. You will ALWAYS be integrating from the top function, and subtracting the bottom function. with U-substitution, you're taking a really hard and complex function, and substituting it with a letter so you can use the FTC to evaluate it.

OpenStudy (jhannybean):

And your welcome :)